# Geometric Optimization and IPA-Induced Dispersion Tuning in Solid-Core Photonic Crystal Fibers

### v1

@@ 1 Introduction
Bandgap guiding in PCFs is based on the photonic bandgap effect, where the photonic crystal structure creates a range of wavelengths that cannot propagate through the crystal. This is achieved by arranging air holes in a regular pattern around the fiber core, allowing for a lower refractive index core. Hollow-core photonic crystal fibers (HC-PCFs) contrast sharply with traditional fibers, where light is guided by a high-index core, resulting in increased loss. In HC-PCFs, the photonic band gap prevents light from propagating in the cladding, effectively trapping and guiding it in the core. This mechanism is beneficial for high-power transmission applications, as it minimizes interaction with the fiber material, reducing loss [20, 21, 22, 23].
@@ 2 Methodology
The PCF designs considered in this study include both hexagonal and circular air-hole geometries. The geometric parameters—core diameter ($d_{c}$), pitch ($\Lambda$), and air-filling ratio ($f$)—were varied systematically to probe their influence on dispersion, attenuation, and nonlinear performance. The air-filling fraction was defined as $f_{c}=2r/\Lambda$ for circular holes and $f_{h}=s/\Lambda$ for hexagonal holes, where $r$ and $s$ denote the radius and side length of the air holes, respectively. To ensure reproducibility and parametric flexibility, the geometry generation was fully automated via scripting.
FDTD simulations employed a mesh resolution of 5 nm and applied perfectly matched layers (PML) with a thickness of 1 $\mu$m in all directions. Modal excitations were introduced using a fundamental Gaussian mode, and convergence was defined by the condition $\Delta n_{\text{eff}}<10^{-6}$ between successive iterations. The FDE analysis used a 10 nm grid step with a conformal meshing algorithm enabled. The simulations were run across the 700–1600 nm spectral range to capture broadband dispersion shifts and mode confinement behavior. The effective mode area was computed from the transverse electric field $\boldsymbol{E}_{t}$ using the expression
@@
[…]
@@
The design of the PCF structure was guided by a benchmark comparison against the Thorlabs NL-2.3-790-02 fiber[34], which has a nominal core diameter of $2.3\pm 0.1\,\mu$m. Although the pitch value in the fabricated fiber is 1.6 $\mu$m, our simulated designs varied $\Lambda$ up to 2.3 $\mu$m for parameter exploration. Two lattice types were considered: PCF-1 with circular air holes and PCF-2 with hexagonal holes. Design parameters were optimized through multiple FDTD and FDE simulations, targeting enhancements in dispersion, confinement, nonlinearity, and loss performance. A schematic view of both PCF types is presented in Figure 1, along with definitions of all geometrical parameters.
@@
Table 1: Calculated physical and optical properties of the proposed PCF2, along with the comparison of experimental results
@@
Finally, we also investigated the impact of contamination by isopropyl alcohol (IPA), modeled by assigning the refractive index of the cladding holes as $n_{\text{IPA}}=1.377$ at 1.55 $\mu$m. The presence of IPA caused a red-shift in the zero-dispersion wavelength and a measurable increase in loss, highlighting the sensitivity of PCFs to ambient contamination. This aspect is particularly relevant for biosensing and supercontinuum applications, where precise control over modal and spectral properties is essential.
@@ 3 Evaluation of Optical Properties
[…]
@@
Diameter $d_{c}$ ($\mu m$) Zero Dispersion Wavelength $\lambda_{0}$ ($nm$)
@@
[…]
@@
$2.4\,\mu$m 22.3 0.04
$2.2\,\mu$m 24.6 0.06
$2.0\,\mu$m 26.9 0.09
$1.8\,\mu$m 28.5 0.13
$1.6\,\mu$m 30.1 0.17
$1.4\,\mu$m 32.7 0.21
Group index $n_{g}$ trends, shown in Figure 4, reveal that higher $n_{g}$ is obtained for smaller cores, particularly at longer wavelengths. At $1.5\,\mu\text{m}$, the group index shifts from $1.51$ for $d_{c}=2.4\,\mu\text{m}$ to $1.57$ for $1.4\,\mu\text{m}$. This has implications for slow-light and dispersion-compensating devices. Moreover, as $d_{c}$ decreases and $n_{g}$ increases, confinement loss and attenuation also rise, introducing practical trade-offs.
@@
To assess contamination sensitivity, we modeled IPA infiltration by adjusting the refractive index of the air holes. As shown in Figure 5, IPA significantly alters the dispersion profile and shifts $\lambda_{0}$. These shifts must be accounted for in sensor design, as IPA-induced refractive index changes can degrade or enhance optical responses depending on the target operating wavelength.
@@ 4 Discussion
[…]
@@ 5 Conclusion
In conclusion, the systematic evaluation and optimization of PCF structures, supported by both [34] and theoretical analyses, enable the development of highly efficient and adaptable photonic devices. The ability to manipulate the optical properties of PCFs through structural modifications and material choices is pivotal for advancing the field of photonics and achieving breakthroughs in various technological applications.
In the context of dispersion engineering in our study by adjusting the core diameter $(d_{c})$, we can precisely tune the ZDW, as shown in Fig. 2. This capability is essential for applications including pulse compression, supercontinuum generation and dispersion compensation in optical communication systems. Also the linear impact of decreasing $(d_{c})$ on the ZDW provides a straightforward approach to achieving the desired dispersion characteristics.
Our analysis, depicted in Fig. 3, indicates that attenuation increases with a decrease in $(d_{c})$. This balance between dispersion management and confinement efficiency highlights the critical need for optimization in PCF design. Balancing low dispersion with minimal attenuation is essential for maximizing the performance of PCFs in various applications. The variations in group index $n_{g}$ with different ($d_{c}$)’s, shown in Fig. 4, are crucial for controlling pulse propagation and group velocity dispersion. This is particularly relevant for high-speed optical communication and precision sensing technologies. Understanding and controlling the group index behavior enables the design of PCFs with tailored propagation characteristics suited to specific applications. The comprehensive dataset provided in Tables 5 and 2 offers valuable insights into the relationships between structural parameters and optical properties. This data is essential for the further refinement and optimization of PCF designs. We can achieve a wide range of optical properties tailored to specific application requirements by systematically varying parameters such as ($d_{c}$) and lattice pitch. Filling the PCF structures with isopropyl alcohol (IPA) significantly modified the dispersion characteristics, as illustrated in Figs. 5(a) and 5(b). This demonstrates the impact of IPA contamination on the performance of PCFs. IPA modifies the refractive index contrast within the fiber, influencing the ZDW and the overall dispersion profile. As a result, this unwanted condition adversely affects the quality of supercontinuum generation. Understanding these effects is crucial for maintaining optimal performance in sensor applications where such contaminants may be present.
(a) Dispersion behavior and ZDW for circular rods with different IPA fillings of PCF-1.
(b) Dispersion behavior and ZDW for circular rods with different IPA fillings of PCF-2.
Figure 5: Dispersion behavior and ZDW for proposed sensor structures.
@@
In summary, the capability to precisely manipulate the optical properties of PCFs through structural and material modifications offers significant potential for advancements in the field of photonics. The continued exploration and optimization of these fibers will drive advancements in optical communication, sensing, and many other technological fields, ensuring that PCFs remain at the forefront of photonic innovation.

### v2

@@ 1 Introduction
Bandgap guiding in PCFs is based on the photonic bandgap effect, where the photonic crystal structure creates a range of wavelengths that cannot propagate through the crystal. This is achieved by arranging air holes in a regular pattern around the fiber core, allowing for a lower refractive index core. Hollow-core photonic crystal fibers (HC-PCFs) differ from traditional fibers in guiding mechanism. However, conventional fibers generally exhibit lower transmission loss in standard telecom windows compared to HC-PCFs [20, 21, 22, 23].
@@ 2 Methodology
The PCF designs considered in this study include both hexagonal and circular air-hole geometries. The geometric parameters—core diameter ($d_{c}$), pitch ($\Lambda$), and air-filling ratio (ff)—were varied systematically to probe their influence on dispersion, attenuation, and nonlinear performance. The air-filling fraction was defined as $f_{c}=2r/\Lambda$ for circular holes and $f_{h}=s/\Lambda$ for hexagonal holes, where $r$ and $s$ denote the radius and side length of the air holes, respectively. To ensure reproducibility and parametric flexibility, the geometry generation was fully automated via scripting.
@@
$A_{\rm eff}=\frac{\left|\iint|\boldsymbol{E}_{t}|^{2}\,dx\,dy\right|^{2}}{\iint|\boldsymbol{E}_{t}|^{4}\,dx\,dy},$
@@
$\eta=\frac{\iint_{\text{core}}S_{z}\,dA}{\iint_{\text{total}}S_{z}\,dA},$
@@
[…]
@@
Table 1: Calculated physical and optical properties of the proposed PCF-2, along with the comparison of experimental results
@@
[…]
@@ 3 Evaluation of Optical Properties
As illustrated in Fig. 2, reducing $d_{c}$ results in a linear blue-shift in $\lambda_{0}$, a behavior that allows for precise tailoring of the dispersion characteristics. The inset highlights the tunability of ZDW across the studied range, which varies from $791\,\text{nm}$ for $d_{c}=2.4\,\mu\text{m}$ to $646\,\text{nm}$ for $d_{c}=1.4\,\mu\text{m}$ (Table 2). This tunability is instrumental for designing PCFs for specific spectral bands.
@@
Diameter $d_{c}$ ($\mu m$) Zero Dispersion Wavelength $\lambda_{0}$ (nmnm)
@@
Fig. 3 shows the associated confinement loss profiles. As expected, smaller core diameters increase $L_{c}$, reflecting the trade-off between tighter confinement and enhanced dispersion control. For example, $L_{c}$ increases from $0.04\,\text{dB/m}$ at $d_{c}=2.4\,\mu\text{m}$ to $0.21\,\text{dB/m}$ at $1.4\,\mu\text{m}$. These values (Table 3) underscore the importance of balancing confinement with other performance metrics.
@@
[…]
Group index $n_{g}$ trends, shown in Fig. 4, reveal that higher $n_{g}$ is obtained for smaller cores, particularly at longer wavelengths. At $1.5\,\mu\text{m}$, the group index shifts from $1.51$ for $d_{c}=2.4\,\mu\text{m}$ to $1.57$ for $1.4\,\mu\text{m}$. This has implications for slow-light and dispersion-compensating devices. Moreover, as $d_{c}$ decreases and $n_{g}$ increases, confinement loss and attenuation also rise, introducing practical trade-offs.
@@
To assess contamination sensitivity, we modeled IPA infiltration by adjusting the refractive index of the air holes. As shown in Fig. 7 and Fig. 8, IPA significantly alters the dispersion profile and shifts $\lambda_{0}$. These shifts must be accounted for in sensor design, as IPA-induced refractive index changes can degrade or enhance optical responses depending on the target operating wavelength.
In line with the properties highlighted in the abstract, we extended our graphical analysis to include the effective mode area ($A_{\text{eff}}$), nonlinear coefficient ($\gamma$), and confinement loss ($L_{c}$). Fig. 5 illustrates the strong inverse relationship between $A_{\text{eff}}$ and $\gamma$, where decreasing $d_{c}$ enhances nonlinear performance by reducing modal area. However, this improvement comes at the expense of higher confinement loss, as depicted in Fig. 6, which shows an exponential rise in $L_{c}$ with decreasing $d_{c}$. These results directly address the reviewer’s concern and confirm the interplay of geometry, nonlinearity, and attenuation. The combined trends emphasize the critical balance required in PCF design: optimizing $d_{c}$ for nonlinear efficiency while maintaining acceptable loss for practical applications such as supercontinuum generation and sensing.
Figure 5: Variation of effective mode area $A_{\text{eff}}$ (left axis, red dashed line with circles) and nonlinear coefficient $\gamma$ (right axis, blue solid line with squares) as a function of core diameter $d_{c}$. The results show that reducing $d_{c}$ decreases $A_{\text{eff}}$ and correspondingly enhances $\gamma$, consistent with the inverse proportionality $\gamma\propto 1/A_{\text{eff}}$.
Figure 6: Confinement loss $L_{c}$ as a function of core diameter $d_{c}$. Smaller core diameters result in significantly higher leakage loss, demonstrating the trade-off between tighter confinement (enhanced $\gamma$) and increased attenuation.
@@ 4 Discussion
A key observation is that reducing the core diameter $d_{c}$ produces a substantial blue-shift in the zero-dispersion wavelength (ZDW). This shift results from enhanced waveguide dispersion, which becomes increasingly dominant over material dispersion as $d_{c}$ is reduced into the sub-micron regime. Our simulations show that $\lambda_{0}$ shifts from $791\,\text{nm}$ to $646\,\text{nm}$ as $d_{c}$ decreases from $2.4\,\mu\text{m}$ to $1.4\,\mu\text{m}$. However, this spectral tunability is accompanied by increases in both attenuation $\alpha$ and confinement loss $L_{c}$, as indicated in Table 3 and Fig. 3. The trade-off between dispersion engineering and modal confinement must therefore be carefully managed.
@@
Clarify the fabrication of solid-core PCFs with hexagonal cladding is relatively straightforward using the stack-and-draw method. However, reducing the core diameter to sub-micron values introduces increased sensitivity to fabrication tolerances and imperfections.
@@
The proposed PCF design is particularly suited for supercontinuum generation (SCG). Its tunable zero-dispersion wavelength (646–791 nm) aligns with common femtosecond pump sources such as Ti:Sapphire ( 800 nm) and Yb-doped fiber lasers ( 1060 nm), while the high nonlinear coefficient ($\gamma$ up to $124\,\text{W}^{-1}\text{km}^{-1}$) enables efficient spectral broadening at low pulse energies. Compared to existing PCFs (e.g., Thorlabs NL series), our design offers the combined advantages of higher nonlinearity and greater robustness against degradation from contaminants such as isopropyl alcohol (IPA), improving both efficiency and practical applicability.
The systematic evaluation and optimization of PCF structures, supported by both [34] and theoretical analyses, enable the development of highly efficient and adaptable photonic devices. The ability to manipulate the optical properties of PCFs through structural modifications and material choices is pivotal for advancing the field of photonics and achieving breakthroughs in various technological applications.
In the context of dispersion engineering in our study by adjusting the core diameter $(d_{c})$, we can precisely tune the ZDW, as shown in Fig. 2. This capability is essential for applications including pulse compression, supercontinuum generation, and dispersion compensation in optical communication systems. Also, the linear impact of decreasing $(d_{c})$ on the ZDW provides a straightforward approach to achieving the desired dispersion characteristics.
Our analysis, depicted in Fig. 3, indicates that attenuation increases with a decrease in $(d_{c})$. This balance between dispersion management and confinement efficiency highlights the critical need for optimization in PCF design. Balancing low dispersion with minimal attenuation is essential for maximizing the performance of PCFs in various applications. The variations in group index $n_{g}$ with different ($d_{c}$)’s, shown in Fig. 4, are crucial for controlling pulse propagation and group velocity dispersion. This is particularly relevant for high-speed optical communication and precision sensing technologies. Understanding and controlling the group index behavior enables the design of PCFs with tailored propagation characteristics suited to specific applications. The comprehensive dataset provided in Tables 5 and 2 offers valuable insights into the relationships between structural parameters and optical properties. This data is essential for the further refinement and optimization of PCF designs. We can achieve a wide range of optical properties tailored to specific application requirements by systematically varying parameters such as ($d_{c}$) and lattice pitch. Filling the PCF structures with isopropyl alcohol (IPA) significantly modified the dispersion characteristics, as illustrated in Figs. 7 and 8. This demonstrates the impact of IPA contamination on the performance of PCFs. IPA modifies the refractive index contrast within the fiber, influencing the ZDW and the overall dispersion profile. As a result, this unwanted condition adversely affects the quality of supercontinuum generation. Understanding these effects is crucial for maintaining optimal performance in sensor applications where such contaminants may be present.
Figure 7: Dispersion behavior and ZDW for circular rods with different IPA fillings of PCF-1.
Figure 8: Dispersion behavior and ZDW for circular rods with different IPA fillings of PCF-2.
@@ 5 Conclusion
In conclusion, tuning the core diameter allows precise control of ZDW $(646\text{--}791~\text{nm})$, enabling applications such as supercontinuum generation. The enhanced nonlinear coefficient $\gamma$ (up to $124~\text{W}^{-1}\text{km}^{-1}$) supports efficient broadband continuum sources. Compared to existing PCFs, the proposed design offers a better balance between dispersion and nonlinearity. Fabrication feasibility is high, though sub-micron cores increase complexity.